OpenStudy (anonymous):

a man is three times as old as his son. in eight years he will be 4 years older than two times his son's age then. how old are they now?

OpenStudy (anonymous):

F age of father now S age of son now

OpenStudy (anonymous):

S= 3S

OpenStudy (anonymous):

F=3S

OpenStudy (anonymous):

F+8 = complete by yourself

OpenStudy (anonymous):

F+8 = 2*(.......)

OpenStudy (anonymous):

F+8 = 2*(.......) + ???

OpenStudy (anonymous):

y=3x Let y be the man's age and x the son's age y+8=2(x+8)+4 y+8=2x+16+4 y+8=2x+20 3x+8=2x+20 x=12 y=36